OpenStudy (anonymous):

how do i go from 2x-4/sqrt of x- x^+4x+3/2x^3/2= 3x^2+4x-3/2x^3/2

OpenStudy (anonymous):

try equation editor, this is not at all clear

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

\[(2x+4)/(\sqrt{x})-(x^2+4x+3)/(2x ^{3/2)}= (3x^2+4x-3)/(2x^3/2)\]

OpenStudy (anonymous):

how do i go from left to right

OpenStudy (anonymous):

looks like you are trying to subtract. one denominator is \[\sqrt{x}\] the other is \[2x^{\frac{3}{2}}=2\sqrt{x^3}\] multiply the first one top and bottom by \[2x\] so that the denominators will be the same.

OpenStudy (anonymous):

you will get \[\frac{2x(x+4)-(x^2+4x+3)}{2\sqrt{x^3}}\]

OpenStudy (anonymous):

last step is to remove parentheses in the numerator and combine like terms

OpenStudy (anonymous):

isnt the sqrtof x = x^1/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{x} \] = x^1/2

OpenStudy (anonymous):

one denominator is \[\sqrt{x}\] the other is \[2\sqrt{x^2}\] if you multiply the first one by \[2x\] you will get \[2x\times \sqrt{x}=2x^{\frac{3}{2}}=2\sqrt{x^2}\]